OpenStudy (anonymous):

Please help! Will give medal! What is the sum of a 7-term geometric series if the first term is -11, the last term is -45,056, and the common ratio is -4? A.) -143,231 B.) -36,047 C.) 144,177 D.) 716,144

OpenStudy (anonymous):

Ok I still don't really get it

OpenStudy (anonymous):

715,144 would be correct

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

just ask me because i am good at math

OpenStudy (anonymous):

Ok @emily.getch what about this one

OpenStudy (anonymous):

that would be 2/3

OpenStudy (anonymous):

ok and how about this?

OpenStudy (anonymous):

it would be A

OpenStudy (anonymous):

told you

OpenStudy (anonymous):

i got to go

OpenStudy (anonymous):

@emily.getch Yayy! Ok last one:

OpenStudy (anonymous):

correction formula for \[Sn=a \frac{ r^n-1 }{ r-1 }\] n=7 a=-11 r=-4 find S7

OpenStudy (anonymous):

\[S7=-11\frac{ \left( -4 \right)^7-1 }{ -4-1 }=-11\frac{ -16384-1 }{ -5 }=\frac{ 11*-16385 }{ 5 }=?\]